# The Labyrinth of Links: Navigating the Associative Maze of Multi-modal LLMs

### v1

@@ Appendix E Comparison with unfiltered Raw Data
Appendix F Ethic Review
As shown in Figure 7, in our human evaluation protocol, we have implemented a comprehensive Ethic Report mechanism to proactively address and mitigate potential ethical concerns. The user interface incorporates dedicated options for participants to report issues related to privacy or other ethical considerations. This is facilitated through a structured reporting system comprising categorical buttons and an open-ended text field for additional context. This approach enables active participant engagement in ethical oversight during the evaluation phase, fostering a collaborative approach to responsible AI development. We prioritize transparency by empowering participants to articulate concerns on potential privacy infringements, algorithmic bias, or instances where the system may induce discomfort or exhibit opaque behavior. This methodology aligns with the best practices delineated by  Zaldivar et al. (2019) Kennedy-Mayo & Gord (2024), which emphasizes the criticality of integrating transparency and user feedback mechanisms to ensure fairness and accountability in machine learning systems.

### v2

@@ Appendix G Comparison with unfiltered Raw Data
Appendix H Ethic Review
As shown in Figure 8, in our human evaluation protocol, we have implemented a comprehensive Ethic Report mechanism to proactively address and mitigate potential ethical concerns. The user interface incorporates dedicated options for participants to report issues related to privacy or other ethical considerations. This is facilitated through a structured reporting system comprising categorical buttons and an open-ended text field for additional context. This approach enables active participant engagement in ethical oversight during the evaluation phase, fostering a collaborative approach to responsible AI development. We prioritize transparency by empowering participants to articulate concerns about potential privacy infringements, algorithmic bias, or instances where the system may induce discomfort or exhibit opaque behavior. This methodology aligns with the best practices delineated by  Zaldivar et al. (2019) Kennedy-Mayo & Gord (2024), which emphasizes the criticality of integrating transparency and user feedback mechanisms to ensure fairness and accountability in machine learning systems.